OpenStudy (raphaelll):

Evaluate the following. Write down the question and all steps of your solution. 16x4+3(2 ³ x5)

zepdrix (zepdrix):

\[\large\rm 16\times4+3\left(2^3\times5\right)\]Remember your PEMDAS or something similar? It tells us how to prioritize these operations. P = Parentheses. That takes the highest priority. So we have to start here: \(\large\rm 2^3\times5\) Now that we're in this room, within the brackets... What do we prioritize first, Exponent, or Multiplication?

OpenStudy (raphaelll):

I don't think i've ever heard of PEMDAS however since the second letter is E i would assume exponent?

zepdrix (zepdrix):

Good. So apply the exponent BEFORE doing multiplication. \(\large\rm 2^3=2\cdot2\cdot2=8\) So within the brackets we have \(\large\rm 8\times5\)

zepdrix (zepdrix):

\[\large\rm 16\times4+3\left(2^3\times5\right)\]\[\large\rm 16\times4+3\left(8\times5\right)\]\[\large\rm 16\times4+3\left(40\right)\]Now that we've done everything in the brackets, we can drop the brackets and just treat it as multiplication.\[\large\rm 16\times4+3\times40\]So what next? Addition or multiplication? :)

OpenStudy (raphaelll):

multiplication first

zepdrix (zepdrix):

k good :d do the multiplication for both of those groups, what do you get?

OpenStudy (raphaelll):

64 + 120

zepdrix (zepdrix):

cool \c:/ looks correct!

zepdrix (zepdrix):

Raphael, hmmmm wasn't that one of the Ninja Turtles? ;O

OpenStudy (raphaelll):

Cool thank you! i really need to remember PEMDAS, made it really simple, and yeah haha my parents named me after the saint though :)

zepdrix (zepdrix):

oh cool :)

OpenStudy (raphaelll):

Thanks for your help!

zepdrix (zepdrix):

I generally prefer to write it like this:Because you are NOT suppose to prioritize multiplication above division. They are on the same level. When you have a problem involving multiplication AND division steps, you'll need to deal with them `from left to right`, you don't necessarily do the multiplication first. Example:\[\large\rm 100\div5\times4\times3\]You would do the \(\large\rm 100\div5\) step FIRST since it is further to the left.